automobile transmission components to auto manufacturers worldwide and is Mopar’s largest supplier. A Trans-Star Engineer has been tasked with evaluating bids for new-generation CNC machinery to be directly linked to the automated manufacturing of high-precision transmission band components. The following are three interest rates that are included in the Vendor bids. Trans-Star will make payments on a semiannual basis only. The engineer is confused about the effective interest rates-what they are annually and over the payment period of 6-months. Answer the following questions:

              Bid #1:   7.25% per year, compounded monthly

              Bid #2:   5.60% per year, compounded quarterly

              Bid #3:   6.25% per quarter, compounded quarterly

  1. Calculate the effective semiannual interest rate for ALL three bids.

b.)    Calculate the effective annual interest rate for ALL three bids.
Answer #1

a) Effective Semi annual interest
Bid #1 = (1+0.725/12)6 - 1 = 3.68%

Bid #2 = (1+0.560/4)2 - 1 = 2.82%

Bid #3 = (1+0.625)2 - 1 = 12.89%

b) Effective annual interest rate
Bid #1 = (1+0.725/12)12 - 1 = 7.50%

Bid #2 = (1+0.560/4)4 - 1 = 5.72%

Bid #3 = (1+0.625)4 - 1 = 27.44%